Solve Second Order (linear and homogenous) Differential Equation 𝑑2𝑦 /𝑑π‘₯2 + 𝑝(π‘₯) 𝑑𝑦/ 𝑑π‘₯ + 𝑄(π‘₯)𝑦 = 0 by writing it in terms of quadratic equation and solving for its roots, mainly the complex roots.Β 


1
Expert's answer
2021-01-18T19:24:03-0500
"\dfrac{d^2y}{dx^2}+p(x)\dfrac{dy}{dx}+Q(x)y=0"

Trivial Solution: For the homogeneous equation above, note that the function "y(x)=0" always satisfies the given equation, regardless what "p(x)" and "Q(x)" are. This constant zero solution is called the trivial solution of such an equation.Β 

The Principle of Superposition: If "y_1" and "y_2" are any two solutions of the homogeneous equation "\dfrac{d^2y}{dx^2}+p(x)\dfrac{dy}{dx}+Q(x)y=0," where "p(x)" and "Q(x)" are continuous on an open interval "I." Then the Wronskian "W(y_1,y_2)(x)" is given byΒ 


"W(y_1,y_2)(x)=Ce^{-\int p(x)dx}"

where "C" is a constant that depends on "y_1" and "y_2," but not on "x."


Then any function of the form "y=C_1y_1+C_2y_2" is also a solution of the equation, for any pair of constants "C_1" and "C_2." Β 


Given a second order linear equation with constant coefficients


"\dfrac{d^2y}{dx^2}+p\dfrac{dy}{dx}+qy=0"

Solve its characteristic equation "r^2+pr+q=0." The general solution depends on the type of roots obtained (use the quadratic formula to find the roots if you are unable to factor the polynomial!):

1. When "p^2-4q>0," there are two distinct real roots "r_1, r_2"


"y=C_1e^{r_1x}+C_2e^{r_2x}"

2. When "p^2-4q<0," there are two complex conjugate roots "r=\lambda\pm\mu i"

"y=C_1e^{\lambda x}\cos\mu x+C_2e^{\lambda x}\sin \mu x"



3. When "p^2-4q=0," there is one repeated real root "r"


"y=C_1e^{rx}+C_2xe^{rx}"

Since "p(x)=p" and "Q(x)=q," being constants, are continuous for every real number, therefore, according to the Existence and Uniqueness Theorem, in each case above there is always a unique solution valid on (βˆ’βˆž, ∞) for any pair of initial conditions "y(x_0)=y_0, y'(x_0)=y'_0."